OpenStudy (anonymous):

-2 1/2 divided by (-3 1/4)

OpenStudy (heisenberg):

First you'll want to change these numbers to 'improper fractions'. Right now, they are a mix of whole numbers and fractions, so to do any operation on them, they need to be in a consistent form. like this: \[-2 * \frac{1}{2} = - \frac{(2*2)+1}{2} = -\frac{5}{2}\]

OpenStudy (anonymous):

-5/2 divided by -13/4 = 10/13

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you welcome........